OpenStudy (anonymous):

evaluate the expression c = 6 x = -2 - c + 9x - 2

OpenStudy (john_es):

\[-c+9x-2=-6+9(-2)-2=-6-18-2=-26\]

OpenStudy (anonymous):

thank you